# Ergodicity, KAM, FPUT

### v1

@@ 2 KAM
The result is that in the coordinates $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}){\buildrel def\over{=}}({\bf A}_{n},{\mbox{\boldmath$\alpha$}}_{n})$, up to a time scale $\sim{\mbox{\boldmath$\omega$}}_{0}^{-1}O(\Lambda^{2}\Big{(}\frac{\varepsilon_{0}}{C}\Big{)}^{-\xi^{n}})$ and in region of phase space of order $\rho_{n}\sim\rho_{0}\Big{(}\Lambda^{2}\frac{\varepsilon_{0}}{C}\Big{)}^{\xi^{n}}$, the motion is quasi periodic motion with spectrum ${\mbox{\boldmath$\omega$}}_{0}$ and Hamiltonian $H_{n}$ with perturbation size $\sim\varepsilon_{n}\sim(\Lambda^{2}\frac{\varepsilon_{0}}{C})^{\xi^{n}}$.
The conclusion is that analyticity in the $\bf A$}} variables is lost: in the limit $n\to\infty$ remains a quasi periodic motion confined for time $\in(0,+\infty)$ on the real (and analytic) torus to which the approximations ${{\bf A}}^{\prime}_{n},{\mbox{\boldmath$\alpha$}}^{\prime}_{n}$ lead.
  • It is also possible to define a sequence of maps $\widetilde{\cal K}_{n}$ defined in the fixed domain ${\mathcal{C}}_{\frac{1}{4}\rho_{0},\frac{1}{4}\kappa_{0}}$ by rescaling the action coordinates of the polydisks by a factor $(\frac{\varepsilon_{n-1}}{C})^{-\frac{1}{2}}=\rho_{n-1}/\rho_{n}$, $n\geq 1$ so that they are all turned into ${\mathcal{C}}_{\frac{1}{4}\rho_{0},\frac{1}{4}\kappa_{0}}$: calling $\widetilde{\eta}_{n}{\buildrel def\over{=}}\frac{\varepsilon_{n-1}}{C}$ the rescaling transformation will rescale time by $\widetilde{\eta}_{n}^{-\frac{1}{2}}$ and will change ${\bf A}_{n}$ into ${\bf A}^{\prime}_{n}=\widetilde{\eta}_{n}^{-\frac{1}{2}}{\bf A}_{n}$ and change the Hamiltonian into:
    $\widetilde{H}_{n}={\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}^{\prime}_{n}+\widetilde{\eta}_{n}^{\frac{1}{2}}\frac{1}{2}({\bf A}^{\prime}_{n}\cdot J_{0}{\bf A}^{\prime}_{n})+\widetilde{\eta}_{n}^{-\frac{1}{2}}f_{n}(\widetilde{\eta}_{n}^{\frac{1}{2}}{\bf A}^{\prime}_{n},{\mbox{\boldmath$\alpha$}}^{\prime}_{n})\,\vtop{\halign{#\cr\rightarrowfill\crcr\kern-1.0pt\nointerlineskip\cr\;${\scriptstyle n\to\infty}$\;\crcr}}\,{\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}^{\prime}_{\infty}$ (2.22)
@@
  • Hence the perturbation $f_{0}$ and the twist $J_{0}$ are, after renormalization, “irrelevant” operators (in Eq.(2.17) both tend to 0 as $n\to\infty$), while the harmonic oscillator ${\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}$ is a “fixed point”: in some sense the transformation has the harmonic oscillator as an attractive fixed point. This completes a proof of the KAM theorem, interpreted in the Renormalization Group frame [4, 20, 21, 22]: it can be classified as a “super-renormalizable” problem, as it requires only a second order perturbation analysis around the trivial fixed point, Eq.(2.7).
@@
In the ${{\bf A}}_{n},{\mbox{\boldmath$\alpha$}}_{n}$ the motion is controlled by a quadratic Hamiltonian ${\mbox{\boldmath$\omega$}}_{0}{{\bf A}}_{n}+\frac{1}{2}{{\bf A}}_{n}J_{n}{{\bf A}}_{n}$ in a portion of action space $O((\frac{\varepsilon_{n}}{C})^{\frac{1}{3}}\rho_{0})$ ”much larger” compared to the previous case, where the size was $O(\frac{\varepsilon_{0}}{C})^{\frac{1}{2}}\rho_{0}$, but $J_{n}=J_{0}$ was simpler.
@@
(3) The analysis of the first iteration is extremely simplified if $f_{0}$ depends only on $\alpha$: this is a Thirring model and in Eq.2.5 it is ${\bf a}=0$, $F_{1}=0$, $F_{2}$ contains only the quadratic term but $F_{3}$ still has to be defined via Eq.2.8 so that at the next iteration dependence on $\bf A$$^{\prime}$ is introduced in $f_{1}$. Nevertheless the above proof remains valid and yields existence of the invariant torus with spectrum ${\mbox{\boldmath$\omega$}}_{0}$ without any condition on the lower bound $J_{0,-}$ as long as it is >0>0. The invariant tori in Thirring models are called “twistless” as no upper bound on $J_{0,-}$ is necessary: for such models a direct proof based on the explicit computation of a series, “Birkhoff series”, converging to the parametric equations of the invariant torus, can be derived, [24] , and covers also some cases in which $J_{0,-}=0$.  [^13] Which do not fit immediately in the above proof: the latter can be adapted to the case $H={\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}+\frac{1}{2}{{\bf A}}^{\perp}\cdot J_{0}{{\bf A}}^{\perp}+f_{0}({{\bf A}}^{\perp},{\mbox{\boldmath$\alpha$}})$ where ${{\bf A}}^{\perp}=(A_{1},A_{2},\ldots,A_{p},0,\ldots 0)$ if $J_{0,j}>0$ for $j\leq p$ and $(J_{0})_{j}=0$ for $j>p$, which generalizes the Thirring models and is even an easy extension of the twistless Hamiltonians theory (as the perturbation may now depend also on $\bf A$$^{\perp}$): the extension is also implied by the analysis in [24] .
@@ 3 FPUT
$H_{0}({\bf p},{\bf q})=H_{0}({\bf P,\bf Q})=\sum_{k=1}^{N-1}\frac{1}{2}(P_{k}^{2}+\omega_{k}^{2}Q_{k})^{2}\,{\buildrel def\over{=}}\,\sum_{k=1}^{N-1}\omega_{k}a_{k}$ (3.2)
where $a_{k}=\frac{(P_{k}^{2}+\omega_{k}^{2}Q_{k}^{2})}{2\omega_{k}}$. The latter is a natural coordinate as the pairs $(a_{k},\varphi_{k})=(a_{k},\arcsin(\frac{\omega_{k}Q_{k}}{\sqrt{2\omega_{k}a_{k}}}))$ are action-angle variables and evolve, in the elastic chain, as $(a_{k},\varphi_{k})\to(a_{k},\varphi_{k}+\omega_{k}t)$; and $({\bf p},{\bf q})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({\bf P,\bf Q})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({\bf a},{\mbox{\boldmath$\varphi$}})$ are symplectic maps.
@@
${{}H_{\varepsilon}({\bf p},{\bf q})=\sum_{k=1}^{N-1}\omega_{k}\alpha_{k}+\varepsilon\sum_{{\bf h}}T_{\bf h}\prod_{i=1}^{4}Q_{h_{i}}={\mbox{\boldmath$\omega$}}_{k}\cdot{\bf a}+\varepsilon f_{0}({\bf a},{\mbox{\boldmath$\varphi$}})}$
@@
${{}f_{0}({\bf a},{\mbox{\boldmath$\varphi$}})=\frac{1}{2^{4}}\sum_{{\bf h}\in Z^{4},{\mbox{\boldmath$\sigma$}}}T_{\bf h}\prod_{j=1}^{4}\sigma_{j}e^{i\sigma_{j}\varphi_{h_{j}}}(\frac{2a_{h_{j}}}{\omega_{j}})^{\frac{1}{2}},}$
where ${\mbox{\boldmath$\sigma$}}=(\sigma_{j})_{j=1}^{4},\sigma_{j}=\pm 1$ and the variables $Q_{k}$ have been expressed in terms of the ${\bf a},{\mbox{\boldmath$\varphi$}}$.
@@
Perturbation theory can be applied to $H_{\varepsilon}$: just look for a canonical transformation generated by a function $\varepsilon\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ as ${\bf a}={\bf a}^{\prime}+\varepsilon{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\varphi$}}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}),{\mbox{\boldmath$\varphi$}}^{\prime}={\mbox{\boldmath$\varphi$}}+\varepsilon{\mbox{\boldmath$\partial$}}_{{\bf a}^{\prime}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ which, if possible, transforms $H_{\varepsilon}({{\bf A}},{\mbox{\boldmath$\varphi$}})$ into $H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ (where $\varphi$ is imagimed expressed as function of $({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$).
$H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})={\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon{\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\varphi$}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}),{\mbox{\boldmath$\varphi$}})+\varepsilon f_{0}({\bf a}^{\prime}+\varepsilon{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\varphi$}}}\Phi,{\mbox{\boldmath$\varphi$}})$ (3.4)
Naturally $\Phi$ would be chosen so that $H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ is ${\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}$. Calculating up to $O(\varepsilon^{2})$ this means that $\Phi$ should be such that $\varepsilon{\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\varphi$}}\Phi({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})+\varepsilon f_{0}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}})$ vanishes. However even this might be impossible if there are ${\mbox{\boldmath$\nu$}}\in Z^{N-1}$ such that ${\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\nu$}}=0$.@
Therefore the best that can be done is to subtract from the above $\Delta_{4}$ the terms which, expressed in the Fourier’s series, have “resonant” harmonics (orthogonal to $\omega$ but $\neq 0$) and define $\Phi$ to cancel, at lowest order, the part of $f_{0}$ left. So $\Phi$ can be chosen so that the new Hamiltonian (once laboriously expressed in the variables $({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$) will contain, up to order $\varepsilon^{2}$, only resonant harmonics (including ${\mbox{\boldmath$\nu$}}=0$).
The new Hamiltonian will have the form $H_{\varepsilon}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})={\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon{\overline{H}}_{1}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})+\varepsilon^{2}H_{2}({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$ where ${\overline{H}}_{1}$ is entirely resonant.
@@
However truncating the sum to an order $s\geq 1$ and neglecting the $\varepsilon^{s+1}H_{\varepsilon,s+1}$ might turn out to be a simpler system. This happens in the present case if $N$ is prime or is not a power of $2$, as pointed out in [17], simply because there are no ${\mbox{\boldmath$\nu$}}\neq 0$ with ${\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}=0$ in the Fourier’s expansion of the potential (i.e. no resonances $\nu$ with $|{\mbox{\boldmath$\nu$}}|\leq 4$). The terms of order $\leq\varepsilon^{2}$ in ${\overline{H}}_{1},{\overline{H}}_{2}$ not only do not depend at all on the angles $\mbox{\boldmath$\varphi$}$$^{\prime}$, but ${\overline{H}}_{1}=0$ and the terms in ${\overline{H}}_{2}$ can be collected in the form $\varepsilon^{2}{\bf a}^{\prime}\cdot J{\bf a}^{\prime}$, where $J$ is explicitly computed in terms of the spectrum ${\mbox{\boldmath$\omega$}}_{0}$ of the elastic chain and furthermore it is $\det J\neq 0$.
Hence the Hamiltonian ${\overline{H}}_{\varepsilon}={\mbox{\boldmath$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon^{2}{\overline{H}}_{2}({\bf a}^{\prime})$ is integrable and, as computed in [17], is found to be:
@@
The surprising simplicity of this normal form, and the $\omega_{h}\neq 0,h=1,\ldots,N-1$, immediately allows to check that ${\mbox{\boldmath$\omega$}}\cdot{\bf a}+\frac{\varepsilon}{2}{\bf a}\cdot J{\bf a}$ has a Jacobian matrix $J$ constant and $\det J\neq 0$.
@@
The spectra of the motions of ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ will have the form ${\mbox{\boldmath$\omega$}}={\mbox{\boldmath$\omega$}}_{0}+\varepsilon J{\bf a}^{\prime}$ and $KAM$ can be applied to perturbations of the considered integrable model. Let $\Omega$ be an open region in phase space, in which the ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ is defined: then $\Omega$ is covered by the tori corresponding to the actions $\bf a$}: e.g. $0<H_{\varepsilon}({\bf a}^{\prime})<E$.
The $\det J\neq 0$ implies that the set in $\Omega$ of invariant tori for whose spectra verify the inequality $|{\mbox{\boldmath$\omega$}}\cdot{\mbox{\boldmath$\nu$}}|>C|{\mbox{\boldmath$\nu$}}|$ fails, for some $0\neq{\mbox{\boldmath$\nu$}}\in Z^{N}$, has volume $C\varepsilon^{-2}\kappa|\Omega|$, for a constant $\kappa>0$: hence the for almost points in $\Omega$ the spectrum $\omega$ verifies a Diophantine inequality.
Hence under a perturbation of ${\overline{H}}_{\varepsilon}$ of size $\eta$ a large part of phase space will remain covered by invariant tori quasi periodically run with $C_{\varepsilon}=C\varepsilon^{2}$-Diophantine frequencies if $C_{\varepsilon}$ is large enough compared to $\eta$. The $\beta$-chain being approximated to order $\eta=\varepsilon^{2}$ by the above integrable system ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ will therefore have a large part of a set $Q$ in phase space covered by $C_{\varepsilon}=C\varepsilon^{2}$-Diophantine tori. In particular this applies to the set $Q$ of energy close to the ground state of the linear chain where $H({\bf p},{\bf q})<E$ (if $E>0$ and $\varepsilon$ is small enough), and the fraction of $Q$ covered by invariant tori will approach $1$ as $\varepsilon\to 0$ at fixed $E>0$. Remark: (1) The statement that KAM implies that at small energy $H_{\varepsilon}({\bf p},{\bf q})<E$ “most phase space points are on invariant tori” requires some care. Describing the system in the coordinates $({\bf a}^{\prime},{\mbox{\boldmath$\varphi$}}^{\prime})$ the set ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ depends on $\varepsilon$: but the application of KAM to ${\overline{H}}\varepsilon({\bf a}^{\prime})$ imposes a condition on the same $\varepsilon$. The remark at the end of Sec.2 is useful here, as it stresses that if the KAM theory is applied to cases in which the Diophantine constant has the form $C\varepsilon^{-1}$ the KAM theorem holds provided the perturbation size $\eta\varepsilon^{-1}$ with $h$ small enough.
(2) About the $\alpha$-model it is possible that, applying perturbation theory to the fput-$\alpha$-chain, an equivalent Hamiltonian is obtained which is an $O(\varepsilon^{2})$ perturbation of an integrable Hamiltonian with $J_{\varepsilon}$ still of $O(\varepsilon)$: but this is not (yet?) proved, [28]. See also the heuristic analysis in Appendix D.

### v2

@@ 2 KAM
The result is that in the coordinates $({\bf A}^{\prime},{\mbox{$\alpha$}}^{\prime}){\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}({\bf A}_{n},{\mbox{$\alpha$}}_{n})$, up to a time scale $\sim{\mbox{$\omega$}}_{0}^{-1}O(\Lambda^{2}\Big(\frac{\varepsilon_{0}}{C}\Big)^{-\xi^{n}})$ and in region of phase space of order $\rho_{n}\sim\rho_{0}\Big(\Lambda^{2}\frac{\varepsilon_{0}}{C}\Big)^{\xi^{n}}$, the motion is quasi periodic motion with spectrum ${\mbox{$\omega$}}_{0}$ and Hamiltonian $H_{n}$ with perturbation size $\sim\varepsilon_{n}\sim(\Lambda^{2}\frac{\varepsilon_{0}}{C})^{\xi^{n}}$.
The conclusion is that analyticity in the $\bf A$}} variables is lost: in the limit $n\to\infty$ remains a quasi periodic motion confined for time $\in(0,+\infty)$ on the real (and analytic) torus to which the approximations ${{\bf A}}^{\prime}_{n},{\mbox{$\alpha$}}^{\prime}_{n}$ lead.
  • It is also possible to define a sequence of maps $\widetilde{\cal K}_{n}$ defined in the fixed domain ${\mathcal{C}}_{\frac{1}{4}\rho_{0},\frac{1}{4}\kappa_{0}}$ by rescaling the action coordinates of the polydisks by a factor $(\frac{\varepsilon_{n-1}}{C})^{-\frac{1}{2}}=\rho_{n-1}/\rho_{n}$, $n\geq 1$ so that they are all turned into ${\mathcal{C}}_{\frac{1}{4}\rho_{0},\frac{1}{4}\kappa_{0}}$: calling $\widetilde{\eta}_{n}{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}\frac{\varepsilon_{n-1}}{C}$ the rescaling transformation will rescale time by $\widetilde{\eta}_{n}^{-\frac{1}{2}}$ and will change ${\bf A}_{n}$ into ${\bf A}^{\prime}_{n}=\widetilde{\eta}_{n}^{-\frac{1}{2}}{\bf A}_{n}$ and change the Hamiltonian into:
    $\widetilde{H}_{n}={\mbox{$\omega$}}_{0}\cdot{\bf A}^{\prime}_{n}+\widetilde{\eta}_{n}^{\frac{1}{2}}\frac{1}{2}({\bf A}^{\prime}_{n}\cdot J_{0}{\bf A}^{\prime}_{n})+\widetilde{\eta}_{n}^{-\frac{1}{2}}f_{n}(\widetilde{\eta}_{n}^{\frac{1}{2}}{\bf A}^{\prime}_{n},{\mbox{$\alpha$}}^{\prime}_{n})\,\vtop{\halign{#\cr\rightarrowfill\crcr\kern-1.0pt\cr\;${\scriptstyle n\to\infty}$\;\crcr}}\,{\mbox{$\omega$}}_{0}\cdot{\bf A}^{\prime}_{\infty}$ (2.22)
@@
  • Hence the perturbation $f_{0}$ and the twist $J_{0}$ are, after renormalization, “irrelevant” operators (in Eq.(2.17) both tend to 0 as $n\to\infty$), while the harmonic oscillator ${\mbox{$\omega$}}_{0}\cdot{{\bf A}}$ is a “fixed point”: in some sense the transformation has the harmonic oscillator as an attractive fixed point. This completes a proof of the KAM theorem, interpreted in the Renormalization Group frame [4, 20, 21, 22]: it can be classified as a “super-renormalizable” problem, as it requires only a second order perturbation analysis around the trivial fixed point, Eq.(2.7).
@@
In the ${{\bf A}}_{n},{\mbox{$\alpha$}}_{n}$ the motion is controlled by a quadratic Hamiltonian ${\mbox{$\omega$}}_{0}{{\bf A}}_{n}+\frac{1}{2}{{\bf A}}_{n}J_{n}{{\bf A}}_{n}$ in a portion of action space $O((\frac{\varepsilon_{n}}{C})^{\frac{1}{3}}\rho_{0})$ ”much larger” compared to the previous case, where the size was $O(\frac{\varepsilon_{0}}{C})^{\frac{1}{2}}\rho_{0}$, but $J_{n}=J_{0}$ was simpler.
@@
(3) The analysis of the first iteration is extremely simplified if $f_{0}$ depends only on $\alpha$: this is a Thirring model and in Eq.2.5 it is ${\bf a}=0$, $F_{1}=0$, $F_{2}$ contains only the quadratic term but $F_{3}$ still has to be defined via Eq.2.8 so that at the next iteration dependence on $\bf A$$^{\prime}$ is introduced in $f_{1}$. Nevertheless the above proof remains valid and yields existence of the invariant torus with spectrum ${\mbox{$\omega$}}_{0}$ without any condition on the lower bound $J_{0,-}$ as long as it is >0>0. The invariant tori in Thirring models are called “twistless” as no upper bound on $J_{0,-}$ is necessary: for such models a direct proof based on the explicit computation of a series, “Birkhoff series”, converging to the parametric equations of the invariant torus, can be derived, [24] , and covers also some cases in which $J_{0,-}=0$.  [^13] Which do not fit immediately in the above proof: the latter can be adapted to the case $H={\mbox{$\omega$}}_{0}\cdot{{\bf A}}+\frac{1}{2}{{\bf A}}^{\perp}\cdot J_{0}{{\bf A}}^{\perp}+f_{0}({{\bf A}}^{\perp},{\mbox{$\alpha$}})$ where ${{\bf A}}^{\perp}=(A_{1},A_{2},\ldots,A_{p},0,\ldots 0)$ if $J_{0,j}>0$ for $j\leq p$ and $(J_{0})_{j}=0$ for $j>p$, which generalizes the Thirring models and is even an easy extension of the twistless Hamiltonians theory (as the perturbation may now depend also on $\bf A$$^{\perp}$): the extension is also implied by the analysis in [24] .
@@ 3 FPUT
$H_{0}({\bf p},{\bf q})=H_{0}({\bf P,\bf Q})=\sum_{k=1}^{N-1}\frac{1}{2}(P_{k}^{2}+\omega_{k}^{2}Q_{k})^{2}\,{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}\,\sum_{k=1}^{N-1}\omega_{k}a_{k}$ (3.2)
where $a_{k}=\frac{(P_{k}^{2}+\omega_{k}^{2}Q_{k}^{2})}{2\omega_{k}}$. The latter is a natural coordinate as the pairs $(a_{k},\varphi_{k})=(a_{k},\arcsin(\frac{\omega_{k}Q_{k}}{\sqrt{2\omega_{k}a_{k}}}))$ are action-angle variables and evolve, in the elastic chain, as $(a_{k},\varphi_{k})\to(a_{k},\varphi_{k}+\omega_{k}t)$; and $({\bf p},{\bf q})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({\bf P,\bf Q})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({\bf a},{\mbox{$\varphi$}})$ are symplectic maps.
@@
${{}H_{\varepsilon}({\bf p},{\bf q})=\sum_{k=1}^{N-1}\omega_{k}\alpha_{k}+\varepsilon\sum_{{\bf h}}T_{\bf h}\prod_{i=1}^{4}Q_{h_{i}}={\mbox{$\omega$}}_{k}\cdot{\bf a}+\varepsilon f_{0}({\bf a},{\mbox{$\varphi$}})}$
@@
${{}f_{0}({\bf a},{\mbox{$\varphi$}})=\frac{1}{2^{4}}\sum_{{\bf h}\in Z^{4},{\mbox{$\sigma$}}}T_{\bf h}\prod_{j=1}^{4}\sigma_{j}e^{i\sigma_{j}\varphi_{h_{j}}}(\frac{2a_{h_{j}}}{\omega_{j}})^{\frac{1}{2}},}$
where ${\mbox{$\sigma$}}=(\sigma_{j})_{j=1}^{4},\sigma_{j}=\pm 1$ and the variables $Q_{k}$ have been expressed in terms of the ${\bf a},{\mbox{$\varphi$}}$.
@@
Perturbation theory can be applied to $H_{\varepsilon}$: just look for a canonical transformation generated by a function $\varepsilon\Phi({\bf a}^{\prime},{\mbox{$\varphi$}})$ as ${\bf a}={\bf a}^{\prime}+\varepsilon{\mbox{$\partial$}}_{{\mbox{$\varphi$}}}\Phi({\bf a}^{\prime},{\mbox{$\varphi$}}),{\mbox{$\varphi$}}^{\prime}={\mbox{$\varphi$}}+\varepsilon{\mbox{$\partial$}}_{{\bf a}^{\prime}}\Phi({\bf a}^{\prime},{\mbox{$\varphi$}})$ which, if possible, transforms $H_{\varepsilon}({{\bf A}},{\mbox{$\varphi$}})$ into $H_{\varepsilon}({\bf a}^{\prime},{\mbox{$\varphi$}})$ (where $\varphi$ is imagimed expressed as function of $({\bf a}^{\prime},{\mbox{$\varphi$}}^{\prime})$).
$H_{\varepsilon}({\bf a}^{\prime},{\mbox{$\varphi$}})={\mbox{$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon{\mbox{$\omega$}}\cdot{\mbox{$\partial$}}_{\mbox{$\varphi$}}\Phi({\bf a}^{\prime},{\mbox{$\varphi$}}),{\mbox{$\varphi$}})+\varepsilon f_{0}({\bf a}^{\prime}+\varepsilon{\mbox{$\partial$}}_{{\mbox{$\varphi$}}}\Phi,{\mbox{$\varphi$}})$ (3.4)
Naturally $\Phi$ would be chosen so that $H_{\varepsilon}({\bf a}^{\prime},{\mbox{$\varphi$}})$ is ${\mbox{$\omega$}}\cdot{\bf a}^{\prime}$. Calculating up to $O(\varepsilon^{2})$ this means that $\Phi$ should be such that $\varepsilon{\mbox{$\omega$}}\cdot{\mbox{$\partial$}}_{\mbox{$\varphi$}}\Phi({\bf a}^{\prime},{\mbox{$\varphi$}})+\varepsilon f_{0}({\bf a}^{\prime},{\mbox{$\varphi$}})$ vanishes. However even this might be impossible if there are ${\mbox{$\nu$}}\in Z^{N-1}$ such that ${\mbox{$\omega$}}\cdot{\mbox{$\nu$}}=0$.
Therefore the best that can be done is to subtract from the above $f_{0}$ the terms which, expressed in the Fourier’s series, have “resonant” harmonics (orthogonal to $\omega$ but $\neq 0$) and define $\Phi$ to cancel, at lowest order, the part of $f_{0}$ left. So $\Phi$ can be chosen so that the new Hamiltonian (once laboriously expressed in the variables $({\bf a}^{\prime},{\mbox{$\varphi$}}^{\prime})$) will contain, up to order $\varepsilon^{2}$, only ’resonant harmonics’ (including ${\mbox{$\nu$}}=0$).
The new Hamiltonian will have the form $H_{\varepsilon}({\bf a}^{\prime},{\mbox{$\varphi$}}^{\prime})={\mbox{$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon{\overline{H}}_{1}({\bf a}^{\prime},{\mbox{$\varphi$}}^{\prime})+\varepsilon^{2}H_{2}({\bf a}^{\prime},{\mbox{$\varphi$}}^{\prime})$ where ${\overline{H}}_{1}$ is entirely resonant.
@@
However truncating the sum to an order $s\geq 1$ and neglecting the $\varepsilon^{s+1}H_{\varepsilon,s+1}$ might turn out to be a simpler system. This happens in the present case if $N$ is prime or is not a power of $2$, as pointed out in [17], simply because there are no ${\mbox{$\nu$}}\neq 0$ with ${\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}=0$ in the Fourier’s expansion of the potential (i.e. no resonances $\nu$ with $|{\mbox{$\nu$}}|\leq 4$). The terms of order $\leq\varepsilon^{2}$ in ${\overline{H}}_{1},{\overline{H}}_{2}$ not only do not depend at all on the angles $\mbox{$\varphi$}$$^{\prime}$, but ${\overline{H}}_{1}=0$ and the terms in ${\overline{H}}_{2}$ can be collected in the form $\varepsilon^{2}{\bf a}^{\prime}\cdot J{\bf a}^{\prime}$, where $J$ is explicitly computed in terms of the spectrum ${\mbox{$\omega$}}_{0}$ of the elastic chain and furthermore it is $\det J\neq 0$.
Hence the Hamiltonian ${\overline{H}}_{\varepsilon}={\mbox{$\omega$}}\cdot{\bf a}^{\prime}+\varepsilon^{2}{\overline{H}}_{2}({\bf a}^{\prime})$ is integrable and, as computed in [17], is found to be:
@@
The surprising simplicity of this normal form, and the $\omega_{h}\neq 0,h=1,\ldots,N-1$, immediately allows to check that ${\mbox{$\omega$}}\cdot{\bf a}+\frac{\varepsilon}{2}{\bf a}\cdot J{\bf a}$ has a Jacobian matrix $J$ constant and $\det J\neq 0$.
@@
The spectra of the motions of ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ will have the form ${\mbox{$\omega$}}={\mbox{$\omega$}}_{0}+\varepsilon J{\bf a}^{\prime}$ and $KAM$ can be applied to perturbations of the considered integrable model. Let $\Omega$ be an open region in phase space, in which the ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ is defined: then $\Omega$ is covered by the tori corresponding to the actions $\bf a$}: e.g. $0<H_{\varepsilon}({\bf a}^{\prime})<E$.
The $\det J\neq 0$ implies that the set in $\Omega$ of invariant tori for whose spectra verify the inequality $|{\mbox{$\omega$}}\cdot{\mbox{$\nu$}}|>C|{\mbox{$\nu$}}|$ fails, for some $0\neq{\mbox{$\nu$}}\in Z^{N}$, has volume $C\varepsilon^{-2}\kappa|\Omega|$, for a constant $\kappa>0$: hence the for almost points in $\Omega$ the spectrum $\omega$ verifies a Diophantine inequality.
Hence under a perturbation of ${\overline{H}}_{\varepsilon}$ of size $\eta$ a large part of phase space will remain covered by invariant tori quasi periodically run with $C_{\varepsilon}=C\varepsilon^{2}$-Diophantine frequencies if $C_{\varepsilon}$ is large enough compared to $\eta$. The $\beta$-chain being approximated to order $\eta=\varepsilon^{2}$ by the above integrable system ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ will therefore have a large part of a set $Q$ in phase space covered by $C_{\varepsilon}=C\varepsilon^{2}$-Diophantine tori. In particular this applies to the set $Q$ of energy close to the ground state of the linear chain where $H({\bf p},{\bf q})<E$ (if $E>0$ and $\varepsilon$ is small enough), and the fraction of $Q$ covered by invariant tori will approach $1$ as $\varepsilon\to 0$ at fixed $E>0$.
Remark: (1) The statement that KAM implies that at small energy $H_{\varepsilon}({\bf p},{\bf q})<E$ “most phase space points are on invariant tori” requires some care. Describing the system in the coordinates $({\bf a}^{\prime},{\mbox{$\varphi$}}^{\prime})$ the set ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ depends on $\varepsilon$: but the application of KAM to ${\overline{H}}_{\varepsilon}({\bf a}^{\prime})$ imposes a condition on the same $\varepsilon$. The remark at the end of Sec.2 is useful here, as it stresses that if the KAM theory is applied to cases in which the Diophantine constant has the form $C\varepsilon^{-1}$ the KAM theorem holds provided the perturbation size $\eta\varepsilon^{-1}$ with $h$ small enough.
(2) About the $\alpha$-model it is possible that, applying perturbation theory to the fput-$\alpha$-chain, an equivalent Hamiltonian is obtained which is an $O(\varepsilon^{2})$ perturbation of an integrable Hamiltonian with $J_{\varepsilon}$ still of $O(\varepsilon)$: but this is not (yet?) proved, [28]. See also the heuristic analysis in Section4.